OpenStudy (anonymous):

Implicit Differentiation. x^3 +y^3 =24xy.

OpenStudy (anonymous):

I ended up with, but its not correct. Not sure what i did wrong. \[y'= \frac{ -3x+24y }{ 3y-24x} \]

OpenStudy (anonymous):

Woops i ignored the squares....

OpenStudy (jhannybean):

\[\large x^3 +y^3 = 24xy\]\[\large \frac{dx}{dy} = 3x^2 + 3y^2 y' = 24(y +xy')\]

OpenStudy (bahrom7893):

3x^2+3y^2(dy/dx)=24(x dy/dx + y)

OpenStudy (jhannybean):

xD

OpenStudy (anonymous):

I caught my own mistake on this one.

OpenStudy (bahrom7893):

aw too slow on my phone

OpenStudy (jhannybean):

Mwuahahaha

OpenStudy (anonymous):

I'm to tired for this, been doing calc all day lol. Sorry. Thanks for the race responses though : )

OpenStudy (jhannybean):

\[\large \frac{dx}{dy} = 3x^2 + 3y^2 y' = 24(y +xy')\]\[\large 3x^2 y' -24xy' = 24y -3x^2 \]\[\large y' = \frac{24y - 3x^2}{3x^2 -24y}\]

OpenStudy (anonymous):

Then divide out 3 from everything to simplify. I got it!!

OpenStudy (jhannybean):

\[\large y' = \frac{24y - 3x^2}{3x^2 -24y}\]\[\large y'=\frac{8y -x^2}{x^2 -8y}\]

OpenStudy (anonymous):

Yup yup!